# ClassInvGen: Class Invariant Synthesis using Large Language Models

### v1

@@ 7. Limitation
ClassInvGen uses generated tests for invariant pruning, but the test suite may include spurious tests that can incorrectly prune valid invariants. The generated tests might not represent valid sequences of method calls; for example, invoking a pop() method before a push() method could fail certain assertions, leading to improper pruning.
@@
Another limitation is the LLM’s context window, which restricts the amount of code that can be processed in a single call. This limitation makes it challenging to handle large codebases. ClassInvGen partially addresses this issue through compositional generation, breaking down the code into manageable parts. Ongoing advancements in LLMs, as highlighted in recent work (Liu et al., 2024; Gao et al., 2023), are also expected to mitigate this limitation.
@@
For future work, we plan to integrate invariant generation with the generation of formal specifications for member functions, enabling LLM a more comprehensive understanding of program behavior. Additionally, we aim to evaluate ClassInvGen on larger and more complex systems beyond Z3, demonstrating its scalability to diverse codebases.
8. Conclusion
@@
In this paper, we describe an approach to leverage LLMs and a lightweight mixed static/dynamic approach to synthesize class invariants. Our experiments on standard C++ data structures as well as a popular and high-assurance codebase demonstrate the feasibility of our approach. Our technique is currently limited by an automated way to integrate the generated tests into the build system of the underlying repo, and the need for developers to validate the invariants. We envision that integrating ClassInvGen with the continuous integration (CI) and pull requests (PR) can aid in scaling the approach to more developers. In future work, we also plan to investigate incorporating developer feedback to repair or strengthen generated invariants.
Acknowledgments
We would like to thank Nikolaj Bjørner for his encouragement and help in inspecting the several generated invariants for the Z3 source code and provide feedback about their utility.
Appendix A Prompt
@@
Figure 22. ClassInvGen Generation system prompt: instruction and task description.
@@
Figure 23. ClassInvGen Generation system prompt: input-output format.
@@
Figure 24. ClassInvGen Generation user prompt template.
@@
Figure 25. ClassInvGen Refinement system prompt: instruction and task description.
@@
Figure 26. ClassInvGen Refinement user prompt template.

### v2

@@ 7 Limitations
At present, ClassInvGen judges an invariant’s correctness with the same test suite that the LLM co-generates alongside that invariant. This design keeps the pipeline fully automated, but it also risks co-adaptation: the model can drift toward invariants that merely fit the behaviours exercised by its own tests, overstating their generality. ClassInvGen uses generated tests for invariant pruning, but the test suite may include spurious tests that can incorrectly prune valid invariants. The generated tests might not represent valid sequences of method calls; for example, invoking a pop() method before a push() method could fail certain assertions, leading to improper pruning.
@@
Another limitation is the LLM’s context window, which restricts the amount of code that can be processed in a single call. This limitation makes it challenging to handle large codebases. ClassInvGen partially addresses this issue through compositional generation, breaking down the code into manageable parts. Ongoing advancements in LLMs, as highlighted in recent work [28, 13], are also expected to mitigate this limitation.
@@
For future work, we plan to integrate invariant generation with the generation of formal specifications for member functions, enabling LLMs to gain a more comprehensive understanding of program behavior. Additionally, we aim to evaluate ClassInvGen on larger and more complex systems beyond Z3, demonstrating its scalability to diverse codebases.
8 Conclusion
@@
In this paper, we describe an approach to leverage LLMs and a lightweight mixed static/dynamic approach to synthesize class invariants. Our experiments on standard C++ data structures as well as a popular and high-assurance codebase demonstrate the feasibility of our approach. Our technique is currently limited by an automated way to integrate the generated tests into the build system of the underlying repo, and the need for developers to validate the invariants. We envision that integrating ClassInvGen with continuous integration (CI) and pull requests (PR) can aid in scaling the approach to more developers. In future work, we also plan to investigate incorporating developer feedback to repair or strengthen generated invariants.
Appendix 0.A Prompt
@@
Figure 14: ClassInvGen Generation system prompt: instruction and task description. Lines 3-7 give the formal definition of a class invariant (must hold before and after every public method).
@@
Figure 15: ClassInvGen Generation system prompt: input-output format.
@@
Figure 16: ClassInvGen Generation user prompt template.
@@
Figure 17: ClassInvGen Refinement system prompt: instruction and task description.
@@
Figure 18: ClassInvGen Refinement user prompt template.
Appendix 0.B Daikon Invariants Frequency Tables
Table 7: Invariants for avl_tree, 11 public methods.
Invariant Count
this->root has only one value 4
this->root._M_t has only one value 4
this->root._M_t.__uniq_ptr_impl<AvlTree::Node, std::default_delete<AvlTree::Node> >._M_t has only one value 4
this[0] has only one value 3
this->n one of { 3, 4 } 3
this[0] != null 2
this->root != null 2
this->root._M_t != null 2
this->root._M_t.__uniq_ptr_impl<AvlTree::Node, std::default_delete<AvlTree::Node> >._M_t != null 2
this->n one of { 1, 2, 3 } 1
this->n one of { 0, 3 } 1
this->n >= 1 1
this->n == 3 1
t.root has only one value 1
t.root._M_t has only one value 1
t.root._M_t.__uniq_ptr_impl<AvlTree::Node, std::default_delete<AvlTree::Node> >._M_t has only one value 1
t.n == 3 1
this->n == return 1
Table 8: Invariants for red_black_tree, 11 public methods.
Invariant Count
this[0] != null 3
this->root != null 3
this->root._M_t != null 3
this->root._M_t.__uniq_ptr_impl<RedBlackTree::Node, std::default_delete<RedBlackTree::Node> >._M_t != null 3
this->n one of { 3, 4, 6 } 3
this->root has only one value 3
this->root._M_t has only one value 3
this->root._M_t.__uniq_ptr_impl<RedBlackTree::Node, std::default_delete<RedBlackTree::Node> >._M_t has only one value 3
this[0] has only one value 2
this->n >= 0 1
(No intersection exists) 1
t.root has only one value 1
t.root._M_t has only one value 1
t.root._M_t.__uniq_ptr_impl<RedBlackTree::Node, std::default_delete<RedBlackTree::Node> >._M_t has only one value 1
t.n == 3 1
No intersection 1
Table 9: Invariants for linked_list, 8 public methods.
Invariant Count
this[0] has only one value 5
this->head has only one value 5
this->head._M_t has only one value 5
this->head._M_t.__uniq_ptr_impl<LinkedList::Node, std::default_delete<LinkedList::Node> >._M_t has only one value 5
this[0] != null 4
this->head != null 4
this->head._M_t != null 4
this->head._M_t.__uniq_ptr_impl<LinkedList::Node, std::default_delete<LinkedList::Node> >._M_t != null 4
this->tail[] elements != null 2
this->tail[].next elements != null 2
this->tail[].next._M_t elements != null 2
this->n >= 0 2
this->n == 0 1
this->tail[].data elements one of { 1 } 1
this->tail[].data one of { [1] } 1
this->n one of { 1 } 1
this->n one of { 1, 2 } 1
this->tail[].data elements >= 1 1
this->tail[].data elements <= this->n 1
this->tail[].data elements one of { 1, 2, 4 } 1
this->tail[].data one of { [1], [2], [4] } 1
this->n one of { 2, 3 } 1
this->tail[].data == [3] 1
this->tail[].data elements == 3 1
Table 10: Invariants for binary_search_tree, 11 public methods.
Invariant Count
this->root has only one value 4
this->root._M_t has only one value 4
this->root._M_t.__uniq_ptr_impl<BinarySearchTree::Node, std::default_delete<BinarySearchTree::Node> >._M_t has only one value 4
this->n one of { 0, 2, 3 } 3
this->n one of { 0, 3 } 3
this[0] has only one value 3
this[0] != null 2
this->root != null 2
this->root._M_t != null 2
this->root._M_t.__uniq_ptr_impl<BinarySearchTree::Node, std::default_delete<BinarySearchTree::Node> >._M_t != null 2
t.root has only one value 1
t.root._M_t has only one value 1
t.root._M_t.__uniq_ptr_impl<BinarySearchTree::Node, std::default_delete<BinarySearchTree::Node> >._M_t has only one value 1
t.n == 3 1
(this->n == return) == (return == orig(this->n)) 1
Table 11: Invariants for heap, 7 public methods.
Invariant Count
this->comp._M_invoker has only one value 7
this->comp._Function_base._M_manager has only one value 7
this[0] has only one value 6
this->data has only one value 6
this->data._Vector_base<int, std::allocator<int> >._M_impl has only one value 6
this->comp has only one value 6
this->comp._Function_base._M_functor has only one value 6
this->comp._Function_base._M_functor._M_unused has only one value 6
this[0] != null 3
this->data != null 3
this->data._Vector_base<int, std::allocator<int> >._M_impl != null 3
this->comp != null 3
this->comp._M_invoker != null 3
this->comp._Function_base._M_functor != null 3
this->comp._Function_base._M_functor._M_unused != null 3
this->comp._Function_base._M_manager != null 3
this->data._Vector_base<int, std::allocator<int> >._M_impl._Vector_impl_data._M_start[] elements >= 1 2
this->data._Vector_base<int, std::allocator<int> >._M_impl._Vector_impl_data._M_start != null 1
this->data._Vector_base<int, std::allocator<int> >._M_impl._Vector_impl_data._M_finish != null 1
this->data._Vector_base<int, std::allocator<int> >._M_impl._Vector_impl_data._M_end_of_storage != null 1
Table 12: Invariants for hash_table, 7 public methods.
Invariant Count
::__digits == "000102…6979899" 3
::__tag == "" 3
this[0] has only one value 3
this->hash_function has only one value 3
this->hash_function._Function_base._M_functor has only one value 3
this->hash_function._Function_base._M_functor._M_unused has only one value 3
this->hash_function._Function_base._M_functor._M_pod_data == "" 3
this->load_factor == 0.75 3
this->table has only one value 3
this->table._Vector_base<… >._M_impl has only one value 2
this->table._Vector_base<… >._M_impl has only one value 1
this->_num_elements one of { 0, 1 } 1
this->_size == 10 1
this->table._Vector_base<… > > > > >._M_impl._Vector_impl_data._M_end_of_storage 1
key._M_dataplus has only one value 1
key._M_dataplus._M_p one of { "key1", "key2" } 1
key._M_string_length == 4 1
this->_num_elements >= 0 1
this->_size >= 0 1
this->_num_elements <= this->_size 1
Table 13: Invariants for vector, 11 public methods.
Invariant Count
this[0] has only one value 9
this->capacity == 5 7
this->data has only one value 5
this->data[] elements one of { 1, 2 } 5
this->n == 2 4
this->data[] == [1, 2] 4
this->n in this->data[] 4
this->data[] == [1] 3
this->n == 0 2
this->n one of { 1 } 2
this->n in return[] 1
return[] == [1, 2] 1
return[] elements one of { 1, 2 } 1
this->capacity == 0 1
this->data == null 1
this->n == 1 1
this->n == v.n 1
this->capacity == v.capacity 1
v.data has only one value 1
v one of { 1, 2 } 1
this->capacity one of { 5 } 1
this->data[] sorted by < 1
this->n <= this->capacity 1
this->n < this->capacity 1
this->n one of { 2, 5 } 1
this->data[] elements == 1 1
this->data[] one of { [1], [1, 2] } 1
return one of { 1, 2 } 1
Table 14: Invariants for queue, 7 public methods.
Invariant Count
this->data has only one value 5
this->maxSize == 10 5
this[0] has only one value 4
this->data[] == [10, 20, 30] 2
this->data[] elements one of { 10, 20, 30 } 2
this->head one of { 0, 1 } 2
this->tail == 3 2
this->n one of { 2, 3 } 2
this->maxSize in this->data[] 2
this[0] != null 2
this->data != null 2
this->data[] elements >= 0 2
this->data[] sorted by < 2
this->head < this->maxSize 2
this->tail < this->maxSize 2
this->data[] elements one of { 1, 2 } 2
this->data[] one of { [1], [1, 2] } 2
this->tail one of { 0, 1, 2 } 2
this->tail in this->data[] 2
this->head - this->tail + this->n == 0 1
this->tail one of { 2, 3, 100 } 1
this->maxSize one of { 10, 160 } 1
this->n < this->maxSize 1
this->head one of { 0, 50 } 1
this->tail >= 0 1
this->n <= this->maxSize 1
this->head <= this->tail 1
this->head one of { 0, 2 } 1
this->n one of { 0, 1 } 1
this->head == other.head 1
this->maxSize == other.maxSize 1
this->data[] == [7, 14] 1
this->data[] elements one of { 7, 14 } 1
this->head == 0 1
other.data has only one value 1
other.tail == 2 1
this->head one of { 0, 1, 2 } 1
this->head - this->tail + return == 0 1
return one of { 0, 1, 2 } 1
Table 15: Invariants for stack, 6 public methods.
Invariant Count
this->maxSize one of { 10, 160, 1280 } 4
this->data[] sorted by < 4
this->data[] elements < this->maxSize 4
this->n < this->maxSize 3
this[0] != null 3
this->data != null 3
this[0] has only one value 2
this->data[] elements >= 0 2
this->n one of { 0, 1, 2 } 1
this->maxSize == other.maxSize 1
this has only one value 1
this->data has only one value 1
this->n == 2 1
this->maxSize == 10 1
other.data has only one value 1
other.data[] == [1, 2] 1
other.data[] elements one of { 1, 2 } 1
this->n <= this->maxSize 1
Appendix 0.C Additional Implementation Details
This appendix provides additional details and examples for the implementation of ClassInvGen.
0.C.1 Code Instrumentation and Invariant Examples
Figure 19 shows how a public method is instrumented with invariant checks, and Figure 20 shows an example of an incorrect invariant.
1bool AvlTree::empty() {
2 check_invariant();
3 auto ret = empty_original();
4 check_invariant();
5 return ret;
6}
7
8bool AvlTree::empty_original() {
9 return n == 0;
10}
(a) AvlTree instrumented with invariants
1void AvlTree::check_invariant() {
2 std::function<bool(const std::unique_ptr<Node>&)>
3 is_balanced = [&](const std::unique_ptr<Node>& node) -> bool {
4 if (!node) return true;
5 int left_height = height(node->left);
6 int right_height = height(node->right);
7 if (std::abs(left_height - right_height) > 1)
8 return false;
9 return is_balanced(node->left) &&
10 is_balanced(node->right);
11 };
12 assert(is_balanced(root));
13}
(b) Example of a correct AvlTree class invariant
Figure 19: AvlTree instrumentation and invariant example
1void AvlTree::check_invariant() {
2 assert(height(root) == get_height(root));
3}
Figure 20: Example of an incorrect invariant of AvlTree because there is no get_height method
0.C.2 Test Generation Examples
Figure 21 shows an example of a generated test suite used for filtering invariants.
1int main() {
2 // Test Case 1: Basic insertions and traversals
3 {
4 AvlTree tree;
5 tree.insert(10); tree.insert(20); tree.insert(5);
6 tree.in_order_traversal();
7 tree.pre_order_traversal();
8 }
9 // Test Case 2: Size, height, empty checks
10 {
11 AvlTree tree;
12 tree.insert(10); tree.insert(20);
13 tree.size(); tree.height(); tree.empty();
14 }
15 // ... more test cases ...
16}
Figure 21: A test suite generated for AvlTree
0.C.3 Refinement Details
ClassInvGen implements a feedback loop for refining failing invariants. Figure 22(a) shows an error message, while Figures 22(b) and 23 show the BST invariant before and after refinement.
1avl_tree.cpp: In lambda function:
2avl_tree.cpp:21:16: error: use of
3’is_balanced’ before deduction of ’auto’
(a) gcc compiler error messages
1void AvlTree::check_invariant() {
2  auto is_bst = [&](const std::unique_ptr<Node>& node,
3 const T& min, const T& max) -> bool {
4 if (!node) return true;
5 if (node->data <= min $|$
6 node->data >= max) return false;
7 return is_bst(node->left, min, node->data) &&
8 is_bst(node->right, node->data, max);
9 };
10 assert(is_bst(root, std::numeric_limits<T>::min(),
11 std::numeric_limits<T>::max()));
12}
(b) BST property before refinement
1void AvlTree::check_invariant() {
2  std::function<bool(const std::unique_ptr<Node>&, const T&, const T&)>
3 is_bst = [&](const std::unique_ptr<Node>& node, const T& min, const T& max) -> bool {
4 if (!node) return true;
5 if (node->data <= min $|$ node->data >= max) return false;
6 return is_bst(node->left, min, node->data) &&
7 is_bst(node->right, node->data, max);
8 };
10}
Figure 23: BST property after refinement: auto is changed to explicit declarations